# Impact of the relativistic Cowling approximation on shear and interface modes of neutron stars

### v1

@@ Abstract
We investigate shear and interface modes excited in neutron stars with an elastic crust in the full general relativistic framework and compare them to the results obtained within the relativistic Cowling approximation. We observe that the Cowling approximation has virtually no impact on the frequencies or the eigenfunctions of the shear modes; in contrast, the interface modes that arise due to the discontinuities of the shear modulus experience a shift in frequency by several percent when applying the Cowling approximation. Furthermore, we extend a scheme based on the properties of Breit-Wigner resonances, which allows us to estimate the damping times of slowly damped modes; our extension can provide an estimation of the damping time even if the features of the Breit-Wigner resonance are incomplete or if some of them violate the underlying linearity assumption. The proposed scheme is also computationally less expensive and numerically more robust and we provide accurate estimates as well as lower bounds for damping times of shear and interface modes.
@@ I Introduction
A precise understanding of the spectrum of neutron stars is, hence, crucial to extracting information from current as well as potential future observations. Their high compactness calls for a general relativistic treatment which considerably complicates the perturbation equations that need to be solved and there are numerous publications approaching this problem, employing different approximations or techniques [32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. One of the most commonly applied approximations is the so-called Cowling approximation [42], in which perturbations of the spacetime are neglected, resulting in a considerable simplification of the equations [43, 44, 45, 46, 47, 48]. The Cowling approximation is well-known for overestimating the frequency of the fundamental ($f$-) mode and the pressure-restored (acoustic) ($p$-) modes [49, 50].
@@
The layout of this paper is as follows. In Sec. II, we show the equations we solve in order to construct equilibrium models and introduce the employed equation of state along with the shear modulus of the elastic crust. In Sec. III, we explain the underlying definitions for the perturbation equations, both in the Cowling approximation as well as in full General Relativity. In the following Sec. IV, we discuss the eigenfrequencies and eigenfunctions of the shear and interface modes and compare their results obtained in the Cowling approximation and in full general Relativity. We also determine values or constraints for their damping times using a method that we detail in App. B. We conclude the study in Sec. V. Unless otherwise mentioned, we use units in which $c=G=1$, where $c$ and $G$ denote the speed of light and the gravitational constant, respectively.
@@ II.2 Equation of State
We employ a zero temperature parametrized EOS based on nuclear parameters that have been proposed by Oyamatsu and Iida [60, 61] (hereafter referred to as OI-EOS). The bulk energy per nucleon is generally given as a function of the baryon number density, $n_{\textrm{b}}$, and an asymmetry parameter, $\tilde{\alpha}$,
@@ III Perturbation Equations
With those ingredients, the perturbation equations concerning the interior of the star can be derived. First, for the full general relativistic case, we show the perturbation equations valid in the elastic crust in Appendix A; the equations (and boundary conditions) relevant for a perfect fluid were first published in Ref. [34] but can also be found in Ref. [56], where the definitions of the perturbation variables is identical to those in the present paper. At the boundaries of the elastic crust, where the shear modulus $\mu$ is discontinuous, we have to impose the continuity of the functions $\tilde{W}$ and $T_{2}:=\delta\Sigma_{r}^{\ \theta}$ as well as the metric functions $H_{0}$, $H_{1}$, and $K$ as junction conditions [70, 56]. In the exterior of the star, we solve the Zerilli equation using the method employing complex coordinates as detailed in Ref. [36]. Second, for the perturbation equations as well as boundary and junction conditions and the corresponding in the Cowling approximation, we refer the reader to Apps. A to C in Ref. [57], where the solution strategy is explained in detail.
IV Results
@@
Having laid out the perturbation equations that need to be solved both in full General Relativity as well as in the Cowling approximation, we now turn to the analysis of the eigenfrequencies within those two formalisms. It is well-known that the frequencies of the fundamental ($f$-) and pressure-restored ($p$-)modes are considerably overestimated within the Cowling approximation [49, 50] (see also Fig. 1); hence, we will focus on the impact of the Cowling approximation on the $s$-modes that arise due to the elasticity of the crust and the two $i$-modes present in the stellar model considered in this study due to the two boundaries (where the shear modulus, $\mu$, discontinuously becomes zero) of the elastic crust, i.e., the boundaries between the envelope and crust surface and between the phase of cylindrical nuclei and the phase of slablike nuclei (which is the crust-core interface). In particular, we will investigate a number of NS models that we construct using a specific OI-EOS described in Sec. II.2 for several different central energy densities. Further, we also construct NS models, where we substitute the OI-EOS with an EOS with a constant speed of sound for a higher-density region as explained in the same section.
IV.1 Shear Modes
@@
$s_{1}$ 1.4241 1.4248 0.05%
$s_{2}$ 2.3354 2.3363 0.04%
$s_{3}$ 3.0247 3.0210 0.12%
$s_{4}$ 3.6367 3.6340 0.07%
$s_{5}$ 4.4687 4.4704 0.04%
$s_{6}$ 5.2418 5.2324 0.18%
In Fig. 1, we show a comparison of the spectra for the $l=2$ modes excited in the NS model constructed with the original OI-EOS with a central energy density of $\epsilon_{c}=0.536\times 10^{15}\textrm{g}/\textrm{cm}^{3}$. In this figure, the vertical axes, $|A_{\rm in}|$ and $\Delta$, denote some kind of error functions given in Appendix B and in Ref. [57]. One can determine the eigenfrequencies with these quantities, i.e., the frequencies with $A_{\rm in}=0$ in the full general relativistic framework (more precisely, when evaluating $A_{\rm in}$ for real-valued frequencies only, sharp minima of this function are relevant) and with $\Delta=0$ for the Cowling approximation correspond to the eigenfrequencies of the stellar model. As mentioned before, the Cowling approximation impacts the $f$- and $p_{1}$-modes as is visible from the discrepancy between the red (full GR) and blue (Cowling) lines. On the other hand, we find that the $s$-modes are essentially unaffected by the Cowling approximation, which are displayed with dashed lines for clarity (in this way, the strongly overlapping spikes of both curves are visible). We also list the frequencies of the first six $s$-modes in Table 1; the relative difference between the two formalisms remains only below $0.25\,\%$, which may be attributed to numerical accuracy as well as marginal differences in the implementations of the two perturbation codes.
@@
IV.1.1 Damping times of shear modes
Shear modes couple very weakly to the spacetime; therefore, they are very weak emitters of gravitational waves and their damping time is very large. The imaginary part of the complex-valued frequency $\omega$ is, hence, very difficult to determine due to numerical truncation errors when iterating for the zero of $A_{\textrm{in}}(\omega)$ in the complex plane. To solve this difficulty, it may be useful to adopt the method derived by Andersson et al. [72], where the damping time is estimated by using the evaluation of $A_{\textrm{in}}(\omega)$ only with purely real-valued frequencies $\omega$, assuming Im$(\omega)=0$. This method utilizes the ratio $\kappa:=A_{\textrm{out}}/A_{\textrm{in}}$ of the outgoing and ingoing wave amplitudes; we detail this method in Appendix B and extend it in order to also be able to yield constraints on the damping time. In our tests, this method yields very accurate values for the damping times of $f$- and $p$-modes (which, in many cases, can also be determined by direct iteration in the complex plane), proving its reliability. As an example, we detail one such test in App. B.1.
Table 3: Data for the $s_{1}$-mode of the NS model with $\epsilon_{c}=0.696\times 10^{15}\textrm{g}/\textrm{cm}^{3}$ and $\alpha=0.6$ for the method detailed in Appendix B. Both poles and zeroes have been determined with high accuracy; the background model is interpolated on more than 60000 grid points. The value for $\alpha_{n}$ is found by locating the minimum of the Breit-Wigner resonance. Applying the estimates using Eqs. (51) and (53) yields a damping time of that mode of $\tau\approx 4.997\times 10^{6}\,$s.
@@ IV.1.1 Damping times of shear modes
$P_{1}/2\pi$ 0.851911403230840
$Z_{1}/2\pi$ 0.851911403262698
$\alpha_{n}/2\pi$ 0.851911403262706
$P_{2}/2\pi$ 0.851911403294545
$Z_{2}/2\pi$ 0.851912423286366
We apply this method exemplarily to the $s_{1}$-mode of the NS model with $\epsilon_{c}=0.696\times 10^{15}\,\textrm{g}/\textrm{cm}^{3}$ and $\alpha=0.6$. Figure 4 shows the Breit-Wigner resonance (i.e., the amplitude $|A_{\textrm{in}}|$) in red, and the phase of the ratio of the amplitudes, ${\rm Im}(\kappa)/{\rm Re}(\kappa)$, in blue. In order to precisely determine the poles and zeroes in ${\rm Im}(\kappa)/{\rm Re}(\kappa)$, which we list in Table 3, a very high frequency resolution of $\Delta\omega\approx 2\pi\times 10^{-9}\,$z is required. Quoting frequencies to 15 significant digits is physically, obviously, meaningless; even a different interpolation method for the tabulated EOS or a different background resolution will most likely impact those values. However, in our analysis, not the absolute values but the differences between those frequencies are relevant and these do require this large number of significant digits. Even though we can determine both poles and both zeroes, the direct evaluation of Eq. (39) yields unreasonable results as the value for $\alpha_{n}$ comes out smaller than $P_{1}$ (but it should lie between $Z_{1}$ and $P_{2}$ as shown in Table 3). This is due to the fact that the zero $Z_{2}$ $(>Z_{1})$ is comparably far away from the actual eigenfrequency and the assumption of linearity of the amplitudes $A_{\textrm{in}}$ and $A_{\textrm{out}}$, on which the method is built, is violated. However, we may still take the minimum of the Breit-Wigner resonance on the real axis, which is located at $f_{\textrm{s}_{1}}=\alpha_{n}/2\pi\approx 851.9\,$Hz.
Trying to avoid using $Z_{2}$ leaves us with three options. First, we can simply use Eq. (40) which requires the knowledge of the triple $(P_{1},P_{2},\alpha_{n})$ and yields $\tau\approx 4.9968\times 10^{6}\,$s. Second, with only the two poles $P_{1}$ and $P_{2}$ (not requiring $\alpha_{n}$), we can apply Eq. (53) and find the interval $4.9966\times 10^{6}\,\textrm{s}<\tau\leq 7.0663\times 10^{6}\,\textrm{s}$. Third, we can use the zero $Z_{1}$ and the pole $P_{2}$ adjacent to the eigenfrequency and plug them into Eq. (51) which yields $4.1401\times 10^{6}\,\textrm{s}<\tau<4.9975\times 10^{6}\,\textrm{s}$. The two intervals have a small overlap region which also contains the value for $\tau$ obtained using Eq. (40); we deduce $\tau\approx 4.997\times 10^{6}\,$s for the $s_{1}$-mode which is roughly two months. These values are obtained using more than 60000 grid points inside the star. With a 10-fold lower number of grid points, the absolute values of the poles and zeroes marginally shift but the estimate for the damping time remains virtually the same.
@@
Table 4: Damping times of first six $s$-modes of the NS model with $\epsilon_{c}=0.696\times 10^{15}\textrm{g}/\textrm{cm}^{3}$ and $\alpha=0.6$. We find that the damping time of the $s_{1}$-mode is the largest and the next five overtones have damping times that are one to two orders of magnitude smaller.
Mode Damping time
$[10^{6}\,\textrm{s}]$
$s_{1}$ 4.9970
$s_{2}$ 0.4920
$s_{3}$ 0.0587
$s_{4}$ 0.2330
$s_{5}$ 0.2460
$s_{6}$ 0.3410
@@
After having laid out in detail the procedure to determine the damping time of a shear ($s_{1}$-) mode, we continue estimating the damping times of its overtones of the same NS model. The course of action is the same for the next five overtones and we show their damping times in Table 4. We observe that the damping times of the overtones first decrease with increasing order up to the third shear mode ($s_{3}$-mode); subsequently, for higher overtones, the damping times increase again. To confirm this behavior of the damping times, we have calculated them using different resolutions of the background model and find the same results.
IV.1.2 Characteristics of $s$-modes for different values of $l$

### v2

@@ Abstract
We investigate shear and interface modes excited in neutron stars with an elastic crust in the full general relativistic framework and compare them to the results obtained within the relativistic Cowling approximation. We observe that the Cowling approximation has virtually no impact on the frequencies or the eigenfunctions of the shear modes; in contrast, the interface modes that arise due to the discontinuities of the shear modulus experience a considerable shift in frequency when applying the Cowling approximation . Furthermore, we extend a scheme based on the properties of the phase of amplitude ratios, which allows us to estimate the damping times of slowly damped modes; our extension can provide an estimation of the damping time even if the features of the amplitude ratio are incomplete or if some of them violate the underlying linearity assumption. The proposed scheme is also computationally less expensive and numerically more robust and we provide accurate estimates as well as lower bounds for damping times of shear and interface modes. We estimate the damping times also via the quadrupole formula and find that it provides good order-of-magnitude estimates.
@@ I Introduction
A precise understanding of the spectrum of neutron stars is, hence, crucial to extracting information from current as well as potential future observations. Their high compactness calls for a general relativistic treatment which considerably complicates the perturbation equations that need to be solved and there are numerous publications approaching this problem, employing different approximations or techniques [32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. One of the most commonly applied approximations is the so-called Cowling approximation, in which perturbations of the spacetime are neglected, resulting in a considerable simplification of the equations [42, 43, 44, 45, 46, 47, 48]. The Cowling approximation is well-known for overestimating the frequency of the fundamental ($f$-) mode and the pressure-restored (acoustic) ($p$-) modes [49, 50].
@@
The layout of this paper is as follows. In Sec. II, we show the equations we solve in order to construct equilibrium models and introduce the employed equation of state along with the shear modulus of the elastic crust. In Sec. III, we explain the underlying definitions for the perturbation equations, both in the Cowling approximation as well as in full General Relativity. We devote Sec. IV to two methods that allow determining the damping times of long-lived modes. In the following Sec. V, we discuss the eigenfrequencies and eigenfunctions of the shear and interface modes and compare their results obtained in the Cowling approximation and in full general Relativity; we also determine values or constraints for their damping times. We conclude the study in Sec. VI. Unless otherwise mentioned, we use units in which $c=G=1$, where $c$ and $G$ denote the speed of light and the gravitational constant, respectively.
@@ II.2 Equation of State
We employ a barotropic, zero temperature parametrized EOS based on nuclear parameters that have been proposed by Oyamatsu and Iida [60, 61] (hereafter referred to as OI-EOS). The barotropicity of the EOS does not cause harm to the study as the oscillation modes under consideration, i.e. shear and interface modes, are virtually unaffected by this assumption. The bulk energy per nucleon is generally given as a function of the baryon number density, $n_{\textrm{b}}$, and an asymmetry parameter, $\tilde{\alpha}$,
@@ III Perturbation Equations
With those ingredients, the perturbation equations concerning the interior of the star can be derived. First, for the full general relativistic case, we show the perturbation equations valid in the elastic crust in Appendix A; the equations (and boundary conditions) relevant for a perfect fluid were first published in Ref. [34] but can also be found in Ref. [56], where the definitions of the perturbation variables is identical to those in the present paper. At the boundaries of the elastic crust, where the shear modulus $\mu$ is discontinuous, we have to impose the continuity of the functions $\tilde{W}$ and $T_{2}:=\delta\Sigma_{r}^{\ \theta}$ as well as the metric functions $H_{0}$, $H_{1}$, and $K$ as junction conditions [70, 56] owing to the continuity of the first and second fundamental forms; additionally, at the interface between spherical and cylindrical nuclei within the crust, where the shear modulus $\mu$ exhibits a small jump, too, we impose continuity of the Lagrangian pressure perturbation, $X$, given in Appendix A. In the exterior of the star, we solve the Zerilli equation using the method employing complex coordinates as detailed in Ref. [36]. Second, for the perturbation equations as well as boundary and junction conditions and the corresponding in the Cowling approximation, we refer the reader to Apps. A to C in Ref. [57], where the solution strategy is explained in detail.
IV Damping Times of Slowly Damped Modes
We perform our calculations of the oscillation modes in the frequency domain; in the full general relativistic case, this approach yields, in principle, the frequency as well as the damping time at the same time. This method works well for $w$-modes [35] and also for the $f$-mode and low-order $p$-modes [33, 36]. However, due to finite numerical accuracy and truncation errors, determining the damping time of very slowly damped modes (for which the imaginary part of the complex-valued frequency is several orders of magnitude smaller than their real part) is computationally challenging and a direct iteration in the complex plane for a zero of $A_{\textrm{in}}(\omega)$ sometimes impossible. The frequency of such modes can be estimated accurately by analysing the amplitude of the ingoing wave $A_{\textrm{in}}(\omega)$ for purely real-valued frequencies $\omega$; in the following, we discuss to methods how estimates for the damping time can be found in these cases.
IV.1 Features of the Phase of the Amplitude Ratio
First, we discuss a method that has been devised by Andersson et al. [71], which utilizes certain features of the ratio $\kappa:=A_{\textrm{out}}/A_{\textrm{in}}$ of the outgoing and ingoing wave amplitudes for purely real-valued frequencies $\omega$. We extend this method in order to also be able to yield constraints on the damping time in the case that not all desired features can be resolved. As this method is rather technical, we show it including our extensions in full detail in App. B, and repeat only the main results here. In our tests, this method yields very accurate values for the damping times of $f$- and $p$-modes (which, in many cases, can also be determined by direct iteration in the complex plane), proving its reliability. As an example, we detail one such test in App. B.1.
The method investigates the quantity $\textrm{Im}(\kappa)/\textrm{Re}(\kappa)$, which is closely linked to the phase of $\kappa$; it can be argued that this quantity features two poles (denoted by $P_{1}$ and $P_{2}$) and two zeroes (denoted by $Z_{1}$ and $Z_{2}$) in the vicinity of an actual eigenfrequency $\omega_{n}$ of the system (cf. Fig. 4, in which two poles and one zero associated with an $s$-mode can be seen). In the optimal case, both poles and both zeroes can be located and the damping time is then given by
$\tau^{-1}$ $=\sqrt{(\alpha_{n}-P_{1})(P_{2}-\alpha_{n})}$ (23)
$=\sqrt{(\alpha_{n}-Z_{1})(Z_{2}-\alpha_{n})},$ (24)
where $\alpha_{n}$ is the frequency of the mode. This can be determined as the minimum of the Breit-Wigner resonance or by the estimate
$\alpha_{n}$ $=\frac{P_{1}P_{2}-Z_{1}Z_{2}}{P_{1}+P_{2}-Z_{1}-Z_{2}}.$ (25)
It is immediately clear that (mathematically) the knowledge of either the two poles or the two zeroes suffices to determine the damping time $\tau$ when $\alpha_{n}$ can be determined accurately without relying on Eq. (25).
In our analyses, we observed that sometimes the location of $\alpha_{n}$ relative to the poles or zeroes could not be accurately enough determined to avoid using Eq. (25); this may happen since locating a minimum can numerically not be as accurate as locating a pole or zero. Hence, it is useful to also be able to provide constraints on the damping time. We find that
$|Z-P|<\tau^{-1}\lesssim 1.2|Z-P|;$ (26)
in this formula, $P$ and $Z$ denote the one pole and the one zero directly adjacent to the mode frequency $\alpha_{n}$. This means that by simply locating the closest pole and zero, we can estimate the damping time to an error of less than 20%. Second, we can derive the estimate
$0.70\delta\lesssim 2\tau^{-1}<\delta,$ (27)
where $\delta$ is the smaller of the differences between the poles and the zeroes, i.e., $\delta:=\min(P_{2}-P_{1},Z_{2}-Z_{1})$. Last, in the case that a mode has such a large damping time that the poles and zeroes cannot be separated from each other at some frequency resolution $\Delta\omega$, then it is
$\tau>2/\Delta\omega.$ (28)
These are, in a compact way, the main results that we are able to extract from the features of the quantity $\textrm{Im}(\kappa)/\textrm{Re}(\kappa)$ and that allow us to either calculate or constrain the damping time $\tau$ of a mode. Nonetheless, we encourage the reader to delve into App. B for a detailed derivation of these expression.
IV.2 Quadrupole formula
Other than using the features of phase of the amplitude ratio, we can also estimate the damping time of slowly damped modes by employing the well-known quadrupole formula [72]; it can then be calculated via
[…]
The energy $E$ of the mode is given by the maximum amplitude of the kinetic energy [74, 73]
[…]
where $\tilde{V}_{l}$ and $\tilde{W}_{l}$ are the eigenfunctions of the tangential and radial displacement of the corresponding oscillation mode which has the angular frequency $\omega_{l}$; this is a full general relativistic expression and is denoted with “R”; its Newtonian counterpart, denoted with “N”, can be obtained by replacing $(\epsilon+p)\rightarrow\rho$ and $e^{\Lambda-\Phi}\rightarrow 1$ in expression (30).
The standard quadrupole formula (denoted “SQF”) is a Newtonian expression and is given by [75]
[…]
where $\delta\rho_{l}$ denotes the eigenfunction of the perturbed rest-mass density; as a full general relativistic formula is not known, we follow the suggestion of Ref. [73] to substitute $\delta\rho_{l}\rightarrow\delta\epsilon_{l}$ (the Eulerian perturbation of the energy density) in Eq. (31), which is then referred to as “RQF”.
[…]
V Results
@@
Having laid out the perturbation equations that need to be solved both in full General Relativity as well as in the Cowling approximation, we now turn to the analysis of the eigenfrequencies within those two formalisms. It is well-known that the frequencies of the fundamental ($f$-) and pressure-restored ($p$-)modes are considerably overestimated within the Cowling approximation [49, 50] (see also Fig. 1); hence, we will focus on the impact of the Cowling approximation on the $s$-modes that arise due to the elasticity of the crust and the two $i$-modes present in the stellar model considered in this study due to the two boundaries (where the shear modulus, $\mu$, discontinuously becomes zero) of the elastic crust, i.e., the boundaries between the envelope and crust surface and between the phase of cylindrical nuclei and the phase of slablike nuclei (which is the crust-core interface).
In particular, we will investigate a number of NS models that we construct using a specific OI-EOS described in Sec. II.2 for several different central energy densities. Further, we also construct NS models, where we substitute the OI-EOS with an EOS with a constant speed of sound for a higher-density region as explained in the same section.
V.1 Shear Modes
@@
$s_{1}$ 1.4241 1.4105 1.0%
$s_{2}$ 2.3354 2.3188 0.7%
$s_{3}$ 3.0247 3.0116 0.4%
$s_{4}$ 3.6367 3.6243 0.3%
$s_{5}$ 4.4687 4.4648 0.1%
$s_{6}$ 5.2418 5.2387 0.1%
In Fig. 1, we show a comparison of the spectra for the $l=2$ modes excited in the NS model constructed with the original OI-EOS with a central energy density of $\epsilon_{c}=0.536\times 10^{15}\textrm{g}/\textrm{cm}^{3}$. In this figure, the vertical axes, $|A_{\rm in}|$ and $\Delta$, denote some kind of error functions given in Appendix B and in Ref. [57]. One can determine the eigenfrequencies with these quantities, i.e., the frequencies with $A_{\rm in}=0$ in the full general relativistic framework (more precisely, when evaluating $A_{\rm in}$ for real-valued frequencies only, sharp minima of this function are relevant) and with $\Delta=0$ for the Cowling approximation correspond to the eigenfrequencies of the stellar model. As mentioned before, the Cowling approximation impacts the $f$- and $p_{1}$-modes as is visible from the discrepancy between the red (full GR) and blue (Cowling) lines. On the other hand, we find that the $s$-modes are largely unaffected by the Cowling approximation, which is displayed with dashed lines for clarity (in this way, the strongly overlapping spikes of both curves are visible). We also list the frequencies of the first six $s$-modes in Table 1; the relative difference between the two formalisms remains only below $1.0\,\%$, which may be attributed to numerical accuracy as well as marginal differences in the implementations of the two perturbation codes.
@@
V.1.1 Damping times of shear modes
Shear modes couple very weakly to the spacetime; therefore, they are very weak emitters of gravitational waves and their damping time is very large. The imaginary part of the complex-valued frequency $\omega$ is, hence, very difficult to determine due to numerical truncation errors when iterating for the zero of $A_{\textrm{in}}(\omega)$ in the complex plane. Hence, we resort to the two methods discussed in Sec. IV in order to determine the damping times of shear modes.
Table 3: Data for the $s_{1}$-mode of the NS model with $\epsilon_{c}=0.696\times 10^{15}\textrm{g}/\textrm{cm}^{3}$ and $\alpha=0.6$ for the method detailed in Appendix B. Both poles and zeroes have been determined with high accuracy; the background model is interpolated on more than 60000 grid points. The value for $\alpha_{n}$ is found by locating the minimum of the Breit-Wigner resonance. Applying the estimates using Eqs. (60) and (62) yields a damping time of that mode of $\tau\approx 3.511\times 10^{6}\,$s.
@@ V.1.1 Damping times of shear modes
$P_{1}/2\pi$ 0.843561278872844
$Z_{1}/2\pi$ 0.843561278918188
$\alpha_{n}/2\pi$ 0.843561278918193
$P_{2}/2\pi$ 0.843561278963515
$Z_{2}/2\pi$ 0.843562814080813
First, we apply the method based on the ratio of the wave amplitudes exemplarily to the $s_{1}$-mode of the NS model with $\epsilon_{c}=0.696\times 10^{15}\,\textrm{g}/\textrm{cm}^{3}$ and $\alpha=0.6$. Figure 4 shows the Breit-Wigner resonance (i.e., the amplitude $|A_{\textrm{in}}|$) in red, and the phase of the ratio of the amplitudes, ${\rm Im}(\kappa)/{\rm Re}(\kappa)$, in blue. In order to precisely determine the poles and zeroes in ${\rm Im}(\kappa)/{\rm Re}(\kappa)$, which we list in Table 3, a very high frequency resolution of $\Delta\omega\approx 2\pi\times 10^{-9}\,$z is required. Quoting frequencies to 15 significant digits is physically, obviously, meaningless; even a different interpolation method for the tabulated EOS or a different background resolution will most likely impact those values. However, in our analysis, not the absolute values but the differences between those frequencies are relevant and these do require this large number of significant digits. Even though we can determine both poles and both zeroes, the direct evaluation of Eq. (48) yields unreasonable results as the value for $\alpha_{n}$ comes out smaller than $P_{1}$ (but it should lie between $Z_{1}$ and $P_{2}$ as shown in Table 3). This is due to the fact that the zero $Z_{2}$ $(>Z_{1})$ is comparably far away from the actual eigenfrequency and the assumption of linearity of the amplitudes $A_{\textrm{in}}$ and $A_{\textrm{out}}$, on which the method is built, is violated. However, we may still take the minimum of the Breit-Wigner resonance on the real axis, which is located at $f_{\textrm{s}_{1}}=\alpha_{n}/2\pi\approx 843.6\,$Hz.
Trying to avoid using $Z_{2}$ leaves us with three options. First, we can simply use Eq. (49) which requires the knowledge of the triple $(P_{1},P_{2},\alpha_{n})$ and yields $\tau\approx 3.5106\times 10^{6}\,$s. Second, with only the two poles $P_{1}$ and $P_{2}$ (not requiring $\alpha_{n}$), we can apply Eq. (62) and find the interval $3.5106\times 10^{6}\,\textrm{s}<\tau\leq 4.9647\times 10^{6}\,\textrm{s}$. Third, we can use the zero $Z_{1}$ and the pole $P_{2}$ adjacent to the eigenfrequency and plug them into Eq. (60) which yields $2.9088\times 10^{6}\,\textrm{s}\leq\tau<3.5113\times 10^{6}\,\textrm{s}$. The two intervals have a small overlap region which also contains the value for $\tau$ obtained using Eq. (49); we deduce $\tau\approx 3.511\times 10^{6}\,$s for the $s_{1}$-mode which is roughly two months. These values are obtained using more than 60000 grid points inside the star. With a 10-fold lower number of grid points, the absolute values of the poles and zeroes marginally shift but the estimate for the damping time remains virtually the same.
@@
Table 4: Damping times of first six $s$-modes (for $l=2$) of the NS model with $\epsilon_{c}=0.696\times 10^{15}\textrm{g}/\textrm{cm}^{3}$ and $\alpha=0.6$. We find that the damping time of the $s_{1}$-mode is the largest and the next five overtones have damping times that are a few orders of magnitude smaller. PAR denotes the damping times estimated via the phase of the amplitude ratio, while N/SQF and R/RQF are estimates employing variants of the quadrupole formula.
Mode Damping time $[10^{3}\,\textrm{s}]$
PAR N/SQF R/RQF
$s_{1}$ 3511.000 2205.000 3533.000
$s_{2}$ 0169.000 0261.000 0409.000
$s_{3}$ $0$.087 $0$.141 $0$.216
$s_{4}$ 0013.400 0034.100 0051.600
$s_{5}$ 0036.000 0145.000 0212.000
$s_{6}$ 0038.600 0152.000 0213.000
@@
After having laid out in detail the procedure to determine the damping time of a shear ($s_{1}$-) mode by analysing its the features of the amplitude ratio and by employing two variants of the quadrupole formula, we continue estimating the damping times of its overtones of the same NS model. The course of action is the same for the next five overtones and we show their damping times in Table 4. We observe that the damping times of the overtones first decrease with increasing order up to the third shear mode ($s_{3}$-mode); subsequently, for higher overtones, the damping times increase again. The damping time of the $s_{3}$-mode is, somewhat surprisingly, so short that we are able to iterate for the complex frequency $\omega$ directly in the complex plane: We find $\tau_{s_{3}}=87.566\,$s while analyzing the amplitude ratio, in very good agreement, yields $\tau_{s_{3}}=87.544\,$s. To confirm this behavior of the damping times, we have calculated them using different resolutions of the background model and find the same results. Furthermore, we note that the two variants of the quadrupole formula systematically underestimate the damping time and the relativistic modification R/RQF results in larger deviations than the Newtonian variant N/SQF. While we believe the estimate via the phase of the amplitude ratio to be accurate to percent level, both quadrupole formula estimates yield at least an order of magnitude estimate for the damping time of the $s$-modes.
V.1.2 Characteristics of $s$-modes for different values of $l$